# Dynamic Strategy Planning for Efficient Question Answering with Large Language Models

### v1

@@ Abstract
Research has shown the effectiveness of reasoning (e.g., Chain-of-Thought), planning (e.g., SelfAsk), and retrieval augmented generation strategies to improve the performance of Large Language Models (LLMs) on various tasks, such as question answering. However, using a single fixed strategy to answer different kinds of questions is suboptimal in performance and inefficient in terms of generated output tokens and performed retrievals. In our work, we propose a novel technique DyPlan, to induce a dynamic strategy selection process in LLMs, to improve performance and reduce costs in question-answering. DyPlan incorporates an initial decision step to select the most suitable strategy conditioned on the input question and guides the LLM’s response generation accordingly. We extend DyPlan to DyPlan-verify, adding an internal verification and correction process to further enrich the generated answer. Experiments on three prominent multi-hop question answering (MHQA) datasets reveal how DyPlan can improve model performance by 7-13% while reducing the cost by 11-32% relative to the best baseline model. Work completed as part of an internship at Meta.
@@ 1 Introduction
To this end, we propose to induce a human-like cognitive ability in LLMs through our novel technique, DyPlan (Dynamic Planning). As illustrated in Figure 1, DyPlan introduces an initial decision step to select the most suitable strategy conditioned on the input question and then guides the LLM’s response generation to use this strategy. To achieve this behavior, we utilize a multi-turn training paradigm, where the LLM is fine-tuned and calibrated by its own generations. DyPlan provides a cost-effective and adaptive solution that efficiently leverages the strengths of various techniques while minimizing computational overhead.
@@
To evaluate the efficacy of our proposed techniques, we benchmark them on three QA datasets - HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and Musique Trivedi et al. (2022). We consider four primary strategies: (1) Direct answering directly, (2) Reason utilizing Chain-of-Thought Wei et al. (2022), (3) Plan leveraging SelfAsk Press et al. (2023), and (4) Retrieval using external knowledge with RAG Lewis et al. (2020). We use the LLaMa3-8B model Dubey et al. (2024) as our base model. We majorly compare against fine-tuned LLMs utilizing fixed strategies along with other ensemble and dynamic thinking baselines. Results reveal that DyPlan reduces costs by 26-32% along with performance gains of 7% averaged across the datasets over the best baseline. DyPlan-verify further improves performance to an average of 12-13% while providing 11-19% cost reductions. Analyses reveal how DyPlan is better calibrated and generalizable and provide insights into its decision-making and verification ability.
In conclusion, we make these contributions: (1) we propose dynamic strategy planning through DyPlan to improve the performance and cost-efficiency of LLMs for QA, (2) we extend DyPlan to DyPlan-verify introducing verification of correctness to further boost model performance, (3) we conduct extensive experiments and analyses using four major strategies on three complex QA datasets to demonstrate DyPlan’s cost-effectiveness and strong performance.
@@ 2.1 Motivation
Dynamic strategy planning can help reduce inference costs by simply selecting lower-cost strategies to answer simpler questions. In our work, we additionally posit that it can improve model performance as strategy selection can act as an ensembling method. We verify this hypothesis through a simple analysis using two strategies of Direct and Reason. Direct prompts the LLM to directly provide the answer to the question while Reason utilizes Chain-of-Thought Wei et al. (2022) to answer step-by-step. Utilizing these strategies, we prompt Llama3-8B-Instruct Dubey et al. (2024), evaluate using F1 score on 1000 samples from 2WikiMultihopQA Ho et al. (2020) dataset, and show their performances as a Venn diagram in Figure 2. Generally, one can assume that adding reasoning should only allow us to further answer harder questions while still being able to answer all questions that Direct strategy answers correctly. However, there is a significant contribution of 7.1% F1 where Reason is incorrect, but Direct is correct (for example, a mistake or hallucination can lead to an incorrect result). We notice similar patterns across datasets and strategies (shown in § B.1). Such patterns throw light on how the choice of an appropriate strategy can also improve model performance.
@@ 2.2 DyPlan Components
The Verification component is optional and only part of our extended technique DyPlan-verify. Intuitively, when we make a decision to choose a certain strategy, like reasoning or retrieval, we cannot be certain it will be successful, as reasoning may fail and retrieval may get irrelevant results. Therefore, DyPlan should have the ability to correct the course as long as we have some more budget before having to present the final answer. This component assesses the validity of the Execution output by prompting the LLM to leverage its self-knowledge and confidence to evaluate the answer’s reasonableness and correctness. To minimize cost, we implement this component by asking the LLM to simply output yes/no, as shown in Figure 3(C).
@@ 2.3 DyPlan Pipeline Flow
DyPlan majorly achieves cost minimization by dynamic decision-making in the Decision phase and restricting generation output space for each component. The base version of DyPlan (also referred to as DyPlan-base) employs a low-cost Decision-Execution pipeline (pink arrow in Figure 3). On the other hand, our extension technique DyPlan-verify utilizes an iterative loop of Decision-Execution-Verification (orange arrow in Figure 3). If verification fails, the pipeline reverts to the Decision component to select an alternative strategy; otherwise, it exits the loop with the execution answer. This iterative loop runs for a preset number of rounds based on the inference budget or until the Decision runs out of usable strategies. Both pipelines are implemented using multi-turn chat, with each component corresponding to a single turn.
@@ 3 Data Creation for Finetuning
To adhere LLMs with the DyPlan pipeline, we fine-tune the LLM on DyPlan-specific data. To ensure zero human annotation cost, we propose automatic data creation utilizing an existing QA dataset $\mathcal{D}$ and a strategy set $\mathcal{S}$ comprising $n$ strategies. We order $\mathcal{S}=[s_{1},\dots,s_{n}]$ by strategy preference, with $s_{1}$ being the most preferred and $s_{n}$ the least preferred (e.g. in order of cost-efficiency/performance). For each strategy $s\in\mathcal{S}$, we prompt the base LLM on all datapoints $d\in\mathcal{D}$ and evaluate the results against the ground truth. This yields two disjoint subsets for each $s$: $D^{s}_{p}$, comprising datapoints where $s$ produced the correct answer, and $D^{s}_{n}$, comprising the remaining datapoints where $s$ failed to produce the correct answer. Utilizing the base LLM (instead of distilling from larger LLMs) ensures a stronger model self-calibration. Using the positive and negative subsets, we create component-specific data for DyPlan (described below) and train the LLM on the combination of all the component data. We conduct training only on the last-turn response for the multi-turn training instances.
@@ 4 Experimentation Details
We evaluate the models on two major dimensions of performance and cost. For performance, we utilize Exact Match (EM) and F1 score evaluated against the ground truth - higher the better. For cost, we consider the number of generated tokens (# T) and number of retrievals (# R) - lower the better. For DyPlan, we report the aggregated cost metrics across the turns.
@@
    Direct: LLM is prompted to directly provide the final answer. This is the cheapest strategy in terms of cost.
@@
Table 2: Performance and cost metrics for two strategy combinations of Direct-Plan-Reason ([ref1] is reference) and Reason-Retrieval ([ref2] is reference). We mark the best and second-best metrics in bold and underline.
@@ 5 Results
We note that external classifiers help reduce cost but don’t improve model performance - demonstrating the difficulty of the task. Dynamic decision-making frameworks like DRAGIN and Ensemble improve performance but are 2-5x more expensive. To this end, DyPlan provides the best balance with an average reduction of 32% token and 26% retrieval cost along with relative performance improvements of 7% EM and 7% F1. DyPlan-verify further improves performance with average relative gains of 13% EM and 12% F1 while reducing the token and retrieval cost by 11% and 19% respectively. In the best case scenario on 2WikiMultihopQA, DyPlan shows 16% performance gains with 52% cost reductions, and DyPlan-verify shows 24% performance gains while reducing costs by 35%. Overall, DyPlan provides strong cost reductions along with decent performance gains.
@@ 5.1 Other strategy combinations
To demonstrate the generalizability of our technique across strategy combinations, we consider two additional combinations of strategies. The first combination - Direct, Plan, Reason - explores the ability of LLMs to utilize only their self-knowledge to answer the question. The second combination - Reason and Retrieval - explores the LLM’s calibration to decide if it requires any external information to answer the question. We show the results for these combinations in Table 2. Similar to the main results, we observe the superior performance of DyPlan and DyPlan-verify in terms of model performance (average relative gains of 6%-10%) as well as cost reduction (average reduction of 13%-20% tokens and 17%-24% retrievals).
@@ 6.1.1 Decision component of DyPlan
We compare the strategy planning distribution of various techniques with the optimal policy for 2WikiMultihopQA in Figure 6. The major difference is the usage of Plan and Reason which are nearly 0% for Fixed/Classifier approaches. On the other hand, DyPlan-base and DyPlan-verify are closer to the optimal distribution. We quantify this proximity of the probability distributions in terms of KL divergence. DyPlan-base and DyPlan-verify achieve low scores of 0.066 and 0.014, respectively, while the classifier baseline has a high divergence score of 0.35. This highlights the better strategy planning and stronger calibration of DyPlan.
@@ 6.1.2 Verification analysis of DyPlan-verify
We study the verification precision, answer rejection rate (datapoints verified as “no"), and the change in strategy distribution pre and post-verification (in terms of KL divergence relative to optimal strategy) to gain a deeper understanding of the impact of verification in DyPlan-verify. We provide these statistics in Table 3. The huge drops in KL-divergence post-verification demonstrate how verification aids better alignment to the optimal policy and, thus, improves model calibration. The low rejection rate ensures the cost doesn’t increase significantly, while the high verification precision underlines the strong utility of the verification step.
@@ 6.2 Generalization Analysis
To assess the generalization of DyPlan, we fine-tune it on combined data from the three benchmark datasets. To ensure a fair comparison, the combined data comprises 20k datapoints (same as individual data) with equal shares from the three datasets. We compare the performance of this combined-data fine-tuned model with the individual-data fine-tuned model in Figure 7 and note how the performances are nearly similar for both models. The cost analysis for the combined data model (Table 17 in § B.4) reveals at-par levels of cost as well. Thus, this study reveals how the gains provided by DyPlan/DyPlan-verify are generalizable and not overfitting to a single dataset.
@@ 7 Related Works
Question-answering has been a prominent task with various popular benchmarking datasets like SQuAD Rajpurkar et al. (2016, 2018), MS MARCO Nguyen et al. (2016), TriviaQA Joshi et al. (2017), and SearchQA Dunn et al. (2017). These datasets are single-hop, i.e., they require simple reasoning to find the answer and are easier to answer. To develop complex reasoning in models, multi-hop question-answering (MHQA) datasets were developed like HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), Compositional Celebrities Press et al. (2023), and Musique Trivedi et al. (2022). To improve MHQA performance, works explored on improving the reasoning capabilities of LLMs using chain-of-thought Wei et al. (2022); Kojima et al. (2022), auto-CoT Zhang et al. (2023), self-consistency Wang et al. (2023), tree-of-thought Yao et al. (2023a). Another line of work focused on planning by question decomposition like SelfAsk Press et al. (2023), ART Paranjape et al. (2023), decomposed prompting Khot et al. (2022) or using explicit planners like ReWOO Xu et al. (2023), LLMCompiler Kim et al. (2024), stepback Zheng et al. (2024b). Works also focus on using external knowledge through retrieval like RAG Lewis et al. (2020), Self-RAG Asai et al. (2024), CRAG Yan et al. (2024) with recent works like IRCoT Trivedi et al. (2023), FLARE Jiang et al. (2023b), SynCheck Wu et al. (2024), and DRAGIN Su et al. (2024) exploring dynamic retrieval.

### v2

@@ Abstract
Research has shown the effectiveness of reasoning (e.g., Chain-of-Thought), planning (e.g., SelfAsk), and retrieval augmented generation strategies to improve the performance of Large Language Models (LLMs) on various tasks, such as question answering. However, using a single fixed strategy to answer different kinds of questions is suboptimal in performance and inefficient in terms of generated output tokens and performed retrievals. In our work, we propose a novel technique DyPlan, to induce a dynamic strategy selection process in LLMs, to improve performance and reduce computational costs in question-answering. DyPlan incorporates an initial decision step to select the most suitable strategy conditioned on the input question and guides the LLM’s response generation accordingly. We extend DyPlan to DyPlan-verify, adding an internal verification and correction process to further enrich the generated answer. Experiments on three prominent multi-hop question answering (MHQA) datasets reveal how DyPlan can improve model performance by 7-13% while reducing the computational cost by 11-32% relative to the best baseline model. Code for this work can be found at https://github.com/facebookresearch/dyplan. Work completed as part of an internship at Meta.
@@ 1 Introduction
To this end, we propose to induce a human-like cognitive ability in LLMs through our novel technique, DyPlan (Dynamic Planning). As illustrated in Figure 1, DyPlan introduces an initial decision step to select the most suitable strategy conditioned on the input question and then guides the LLM’s response generation to use this strategy. To achieve this behavior, we utilize a multi-turn training paradigm, where the LLM is fine-tuned and calibrated by its own generations. DyPlan provides a computationally cost-effective and adaptive solution that efficiently leverages the strengths of various techniques while minimizing computational overhead.
@@
To evaluate the efficacy of our proposed techniques, we benchmark them on three QA datasets - HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and Musique Trivedi et al. (2022). We consider four primary strategies: (1) Direct answering directly, (2) Reason utilizing Chain-of-Thought Wei et al. (2022), (3) Plan leveraging SelfAsk Press et al. (2023), and (4) Retrieval using external knowledge with RAG Lewis et al. (2020). We use the LLaMa3-8B model Dubey et al. (2024) as our base model. We majorly compare against fine-tuned LLMs utilizing fixed strategies along with other ensemble and dynamic thinking baselines. Results reveal that DyPlan reduces computational costs by 26-32% along with performance gains of 7% averaged across the datasets over the best baseline. DyPlan-verify further improves performance to an average of 12-13% while providing 11-19% computational cost reductions. Analyses reveal how DyPlan is better calibrated and generalizable and provide insights into its decision-making and verification ability.
In conclusion, we make these contributions: (1) we propose dynamic strategy planning through DyPlan to improve the performance and computational cost-efficiency of LLMs for QA, (2) we extend DyPlan to DyPlan-verify introducing verification of correctness to further boost model performance, (3) we conduct extensive experiments and analyses using four major strategies on three complex QA datasets to demonstrate DyPlan’s cost-effectiveness and strong performance.
@@ 2.1 Motivation
Dynamic strategy planning can help reduce inference computational costs by simply selecting lower-cost strategies to answer simpler questions. In our work, we additionally posit that it can improve model performance as strategy selection can act as an ensembling method. We verify this hypothesis through a simple analysis using two strategies of Direct and Reason. Direct prompts the LLM to directly provide the answer to the question while Reason utilizes Chain-of-Thought Wei et al. (2022) to answer step-by-step. Utilizing these strategies, we prompt Llama3-8B-Instruct Dubey et al. (2024), evaluate using F1 score on 1000 samples from 2WikiMultihopQA Ho et al. (2020) dataset, and show their performances as a Venn diagram in Figure 2. Generally, one can assume that adding reasoning should only allow us to further answer harder questions while still being able to answer all questions that Direct strategy answers correctly. However, there is a significant contribution of 7.1% F1 where Reason is incorrect, but Direct is correct (for example, a mistake or hallucination can lead to an incorrect result). We provide a qualitative example in Table 1 highlighting this behavior (more examples in Table 24). We notice similar patterns across datasets and strategies (shown in § B.1). Such patterns shed light on how the choice of an appropriate strategy can also improve model performance.
Question: Which band, Letters to Cleo or Screaming Trees, had more members?
Gold Answer: Letters to Cleo
Direct: Final answer: "Letters to Cleo"
Reason: Letters to Cleo had 4 members: Brian Desveaux, Michael Hubbard, Ken Ray, and Stacy Jones. Screaming Trees had 4 members: Mark Lanegan, Gary Lee Conner, Van Conner, and Barrett Martin. Both bands had the same number of members, 4. Final answer: "same"
Table 1: Qualitative example highlighting how LLMs can yield the right answer by directly answering the question, while reasoning can lead to incorrect answers.
@@ 2.2 DyPlan Components
The Verification component is optional and only part of our extended technique DyPlan-verify. Intuitively, when we make a decision to choose a certain strategy, like reasoning or retrieval, we cannot be certain it will be successful, as reasoning may fail and retrieval may get irrelevant results. Therefore, DyPlan should have the ability to correct the course as long as we have some more budget before having to present the final answer. This component assesses the validity of the Execution output by prompting the LLM to leverage its self-knowledge and confidence to evaluate the answer’s reasonableness and correctness. To minimize computational cost, we implement this component by asking the LLM to simply output yes/no, as shown in Figure 3(C).
@@ 2.3 DyPlan Pipeline Flow
DyPlan majorly achieves computational cost minimization by dynamic decision-making in the Decision phase and restricting generation output space for each component. The base version of DyPlan (also referred to as DyPlan-base) employs a low-cost Decision-Execution pipeline (pink arrow in Figure 3). On the other hand, our extension technique DyPlan-verify utilizes an iterative loop of Decision-Execution-Verification (orange arrow in Figure 3). If verification fails, the pipeline reverts to the Decision component to select an alternative strategy; otherwise, it exits the loop with the execution answer. This iterative loop runs for a preset number of rounds based on the inference budget or until the Decision runs out of usable strategies. Both pipelines are implemented using multi-turn chat, with each component corresponding to a single turn.
@@ 3 Data Creation for Finetuning
To adhere LLMs with the DyPlan pipeline, we fine-tune the LLM on DyPlan-specific data. To ensure zero human annotation cost, we propose automatic data creation utilizing an existing QA dataset $\mathcal{D}$ and a strategy set $\mathcal{S}$ comprising $n$ strategies. We order $\mathcal{S}=[s_{1},\dots,s_{n}]$ by strategy preference, with $s_{1}$ being the most preferred and $s_{n}$ the least preferred (e.g. in order of computational cost-efficiency/performance). For each strategy $s\in\mathcal{S}$, we prompt the base LLM on all datapoints $d\in\mathcal{D}$ and evaluate the results against the ground truth. This yields two disjoint subsets for each $s$: $D^{s}_{p}$, comprising datapoints where $s$ produced the correct answer, and $D^{s}_{n}$, comprising the remaining datapoints where $s$ failed to produce the correct answer. Utilizing the base LLM (instead of distilling from larger LLMs) ensures a stronger model self-calibration. Using the positive and negative subsets, we create component-specific data for DyPlan (described below) and train the LLM on the combination of all the component data. We conduct training only on the last-turn response for the multi-turn training instances.
@@ 4 Experimentation Details
We evaluate the models on two major dimensions of performance and computational cost. For performance, we utilize Exact Match (EM) and F1 score evaluated against the ground truth - higher the better. For computational cost, we consider the number of generated tokens (# T) and number of retrievals (# R) - lower the better. For DyPlan, we report the aggregated cost metrics across the turns.
@@
    Direct: LLM is prompted to directly provide the final answer. This is the cheapest strategy in terms of computational cost.
@@
Table 3: Performance and computational cost metrics for two strategy combinations of Direct-Plan-Reason ([ref1] is reference) and Reason-Retrieval ([ref2] is reference). We mark the best and second-best metrics in bold and underline.
@@ 5 Results
We note that external classifiers help reduce computational cost but don’t improve model performance - demonstrating the difficulty of the task. Dynamic decision-making frameworks like DRAGIN and Ensemble improve performance but are 2-5x more expensive. To this end, DyPlan provides the best balance with an average reduction of 32% token and 26% retrieval cost along with relative performance improvements of 7% EM and 7% F1. DyPlan-verify further improves performance with average relative gains of 13% EM and 12% F1 while reducing the token and retrieval cost by 11% and 19% respectively. In the best case scenario on 2WikiMultihopQA, DyPlan shows 16% performance gains with 52% computational cost reductions, and DyPlan-verify shows 24% performance gains while reducing costs by 35%. Overall, DyPlan provides strong computational cost reductions along with decent performance gains.
@@ 5.1 Other strategy combinations
To demonstrate the generalizability of our technique across strategy combinations, we consider two additional combinations of strategies. The first combination - Direct, Plan, Reason - explores the ability of LLMs to utilize only their self-knowledge to answer the question. The second combination - Reason and Retrieval - explores the LLM’s calibration to decide if it requires any external information to answer the question. We show the results for these combinations in Table 3. Similar to the main results, we observe the superior performance of DyPlan and DyPlan-verify in terms of model performance (average relative gains of 6%-10%) as well as computational cost reduction (average reduction of 13%-20% tokens and 17%-24% retrievals).
@@ 6.1.1 Decision component of DyPlan
We compare the strategy planning distribution of various techniques with the optimal policy for 2WikiMultihopQA in Figure 6. The major difference is the usage of Plan and Reason which are nearly 0% for Fixed/Classifier approaches. On the other hand, DyPlan-base and DyPlan-verify are closer to the optimal distribution. We quantify this proximity of the probability distributions in terms of KL divergence. DyPlan-base and DyPlan-verify achieve low scores of 0.066 and 0.014, respectively, while the classifier baseline has a high divergence score of 0.35. Finally, for a stronger sanity check, we compute the accuracies of the strategy choice (relative to optimal policy) of DyPlan in Table 4. The high improvements relative to other baselines highlight the better strategy planning and stronger calibration of DyPlan, while throwing light towards further possible improvements.
Technique Accuracy
Random 25.8%
Majority 30.8%
Classifier 48.1%
DyPlan 61.7%
Table 4: Accuracy of the Decision component of DyPlan in choosing the right strategy evaluated using the optimal policy on the HotpotQA dataset.
@@ 6.1.2 Verification analysis of DyPlan-verify
We study the verification precision, answer rejection rate (datapoints verified as “no"), and the change in strategy distribution pre and post-verification (in terms of KL divergence relative to optimal strategy) to gain a deeper understanding of the impact of verification in DyPlan-verify. We provide these statistics in Table 5. The huge drops in KL-divergence post-verification demonstrate how verification aids better alignment to the optimal policy and, thus, improves model calibration. The low rejection rate ensures the computational cost doesn’t increase significantly, while the high verification precision underlines the strong utility of the verification step.
@@ 6.2 Generalization Analysis
To assess the generalization of DyPlan, we fine-tune it on combined data from the three benchmark datasets. To ensure a fair comparison, the combined data comprises 20k datapoints (same as individual data) with equal shares from the three datasets. We compare the performance of this combined-data fine-tuned model with the individual-data fine-tuned model in Figure 7 and note how the performances are nearly similar for both models. The cost analysis for the combined data model (Table 20 in § B.4) reveals at-par levels of computational costs as well. Thus, this study reveals how the gains provided by DyPlan/DyPlan-verify are generalizable and not overfitting to a single dataset.
Strategy Ordering EM F1 # T # R
DyPlan
Direct, Plan, Reason, Retrieve 36.1 47.6 42 0.76
Direct, Reason, Plan, Retrieve 35.4 46.8 40 0.69
Reason, Direct, Plan, Retrieve 36.1 47.7 42 0.72
DyPlan-verify
Direct, Plan, Reason, Retrieve 36.7 48.5 53 0.78
Direct, Reason, Plan, Retrieve 36.5 48.5 51 0.72
Reason, Direct, Plan, Retrieve 37.2 48.5 58 0.76
Table 8: Impact of different strategy orderings for DyPlan on downstream performance and computational costs on the HotpotQA dataset.
6.3 Analyzing the Order of Strategies
In this analysis, we study the impact of different strategy ordering $\mathcal{S}$ for DyPlan. Specifically, we consider three different orderings for HotpotQA and show the results in Table 8. Differently ordering the strategies (e.g., Direct, Reason, Plan, Retrieve) can help further reduce the computational cost by 4-7%, but it can also reduce the performance by 1-2%. Similarly, the performance can be optimized by a different ordering (e.g., Reason, Direct, Plan, Retrieve), with improvements upto 1-2% while incurring additional computational costs upto 10%. This puts focus on exploring the choice of the right strategy ordering as a key component for optimization, but we will keep that for future works.
@@ 7 Related Works
Question-answering is a popular task, with wide-spread applications such as document parsing Suvarna et al. (2024), information extraction Parekh et al. (2024c), chatbots Chalkidis et al. (2022); Singhal et al. (2023), summarization Fabbri et al. (2022), as well as great multilingual Parekh et al. (2024a, b) and multimodal Singh et al. (2019); Talmor et al. (2021) applications. Some prominent benchmarking QA datasets include SQuAD Rajpurkar et al. (2016, 2018), MS MARCO Nguyen et al. (2016), TriviaQA Joshi et al. (2017), and SearchQA Dunn et al. (2017). These datasets are single-hop, i.e., they require simple reasoning to find the answer and are easier to answer. To develop complex reasoning in models, multi-hop question-answering (MHQA) datasets were developed like HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), Compositional Celebrities Press et al. (2023), and Musique Trivedi et al. (2022). To improve MHQA performance, works explored on improving the reasoning capabilities of LLMs using chain-of-thought Wei et al. (2022); Kojima et al. (2022), auto-CoT Zhang et al. (2023), self-consistency Wang et al. (2023), tree-of-thought Yao et al. (2023a). Another line of work focused on planning by question decomposition like SelfAsk Press et al. (2023), ART Paranjape et al. (2023), decomposed prompting Khot et al. (2022) or using explicit planners like ReWOO Xu et al. (2023), LLMCompiler Kim et al. (2024), stepback Zheng et al. (2024b). Works also focus on using external knowledge through retrieval like RAG Lewis et al. (2020), Self-RAG Asai et al. (2024), CRAG Yan et al. (2024) with recent works like IRCoT Trivedi et al. (2023), FLARE Jiang et al. (2023b), SynCheck Wu et al. (2024), and DRAGIN Su et al. (2024) exploring dynamic retrieval. The closest approach to our work is Adaptive RAG Jeong et al. (2024) which utilizes a classifier to determine the question complexity and adaptively utilize RAG. In comparison, our work is more generalized to adapt to any kind of prompt/tool and induces deeper thinking in the LLM itself, while being highly cost-effective at the same time.